A 2055 kg elevator accelerates upward at 1.500 m/s? What is the tension in the cable that supports the elevator? KN 0
Answer #1

2055 kg a = 1.50 m/s² up word T a M my ma T-ng T mar ml gra) T = 2055 (9.81+1.50) T = 23 24 2.05 N T 23. 242 x 103 T = 23. 24